OpenStudy (anonymous):

can you help me please someone?!! Point X is located at (2, 4), and Y is located at (−6, 10). Find the y value of the point that lies halfway between X and Y. 2 7 -3 -8

OpenStudy (anonymous):

The midpoint between two points \((x_1, y_1)\) and \((x_2, y_2)\) is given by: \(\huge Midpoint (x, y) = ((\frac{x_1 + x_2}{2}), (\frac{y_1 + y_2}{2}))\)

OpenStudy (anonymous):

yes and after you do that you get (2,7)

OpenStudy (anonymous):

after that i get confused

OpenStudy (anonymous):

How do you get 2???

OpenStudy (anonymous):

\[\frac{2-4}{2} =??\]

OpenStudy (anonymous):

See, a point contains x and y coordinates, the question is asking you for y coordinates of the midpoint that you will calculate...

OpenStudy (anonymous):

lol i didnt read the question right.... one of those days

OpenStudy (anonymous):

You must get your midpoint as (-2, 7).. And now you know out of -2 and 7 which is your y coordinate???

OpenStudy (anonymous):

Getting now???

OpenStudy (anonymous):

So, what will you get as your y coordinate???

OpenStudy (anonymous):

lol 7 sorry i been having one of those off days

OpenStudy (anonymous):

It is okay.. Keep trying and practising more and more..

OpenStudy (anonymous):

ty lol